OpenStudy (anonymous):

If point A is located at (0, –8) and B is located at (4,–5), what is the distance between them?

OpenStudy (anonymous):

a^2+b^2=c^2 Do you know this formula? :-)

OpenStudy (anonymous):

Yes

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) does this help ??

OpenStudy (anonymous):

um ya but i dont understand that forumula like how do you use it